# A Retrieval-Augmented Generation Approach to Extracting Algorithmic Logic from Neural Networks

### v1

@@ Abstract
Reusing existing neural-network components is central to research efficiency, yet discovering, extracting, and validating such modules across thousands of open-source repositories remains difficult. We introduce NN-RAG, a retrieval-augmented generation system that converts large, heterogeneous PyTorch codebases into a searchable and executable library of validated neural modules. Unlike conventional code search or clone-detection tools, NN-RAG performs scope-aware dependency resolution, import-preserving reconstruction, and validator-gated promotion—ensuring that every retrieved block is scope-closed, compilable, and runnable. Applied to 19 major repositories, the pipeline extracted 1,289 candidate blocks, validated 941 (73.0%), and demonstrated that over 80% are structurally unique. Through multi-level de-duplication (exact, lexical, structural), we find that NN-RAG contributes the overwhelming majority of unique architectures to the LEMUR dataset, supplying $\approx$72% of all novel network structures. Beyond quantity, NN-RAG uniquely enables cross-repository migration of architectural patterns—automatically identifying reusable modules in one project and regenerating them, dependency-complete, in another context. To our knowledge, no other open-source system provides this capability at scale. The framework’s neutral specifications further allow optional integration with language models for synthesis or dataset registration without redistributing third-party code. Overall, NN-RAG transforms fragmented vision code into a reproducible, provenance-tracked substrate for algorithmic discovery, offering a first open-source solution that both quantifies and expands the diversity of executable neural architectures across repositories.
@@ 1 Introduction
We address this gap with NN-RAG—a retrieval-first system that turns heterogeneous Python sources into dependency-closed, executable modules accompanied by provenance. The premise is simple: if we can reliably retrieve, assemble, and validate underused design ideas at the right granularity, then we not only accelerate reuse for ablations and baselines but also surface combinations that the literature rarely co-deploys. In parallel, we emit neutral specifications for each module so that a language model can be fine-tuned to internalize these ideas without memorizing upstream implementations.
The goal of this paper is not to pursue leaderboard performance, but rather the discovery and extraction of new ideas. Nevertheless, while introducing new and unique ideas to the LEMUR dataset [lemur_arxiv], the pipeline surfaced a model that currently attains the best accuracy on that benchmark (Fig. 7). This outcome supports our goal: grounding design decisions in retrieved, validator-gated modules can improve both research velocity and model quality. To understand why, we complement headline numbers with targeted ablations, isolating which ingredients contribute most to the observed gains.
@@
Using a companion curation tool (NN-DUP [nndup_repo]) with exact, MinHash/LSH near-deduplication, AST-fingerprint structural deduplication, and a diversity top-up, we find that the overwhelming majority of unique architectures in LEMUR originate from NN-RAG extractions. [lemur_arxiv, nnrag_repo, bigcode_dedup]
@@
Roadmap. Section 2 motivates the retrieval-first stance and summarizes findings; Section 3 explains the related work; Section 4 details the pipeline; Section 5 reports extraction/validation statistics and ablations; Section 6 dicusses the results; Section 7 concludes.
@@ 2 Motivation
Deep learning reuse ought to be easy: the community has already produced a wealth of layers, losses, and architectural ideas; yet those ideas are scattered across heterogeneous PyTorch repositories with incompatible conventions and hidden dependencies. In practice, porting even a small idea requires three costly steps—discovery at the right granularity, closing transitive imports without breakage, and validating an artifact that reviewers can actually run. This friction slows iteration, inhibits fair comparisons, and blunts the scientific value of open code.
@@
Our stance in NN-RAG is therefore retrieval first. Rather than ask a language model to synthesize code from scratch, we ground assembly in retrieved, verifiable sources, and only optionally invoke generation downstream from neutral specifications. Retrieval-augmented methods are known to improve factual grounding by supplementing a model’s parametric memory with explicit context; in our setting, that means the system privileges concrete, inspectable code over unconstrained paraphrase [Lewis2020RAG, HallucinationSurvey2023]. This design choice directly serves our aims: it turns the open-source corpus into a dependable stream of reusable ideas while curbing the brittleness and hallucination risks of pure free-form generation.
@@
Reliability and safety concerns reinforce this choice. Controlled studies show that AI coding assistance can encourage plausible but insecure solutions when specifications are implicit [Perry2023Insecure]. NN-RAG counters this with validator-gated artifacts: import-preserving regeneration followed by static/dynamic checks, so only dependency-complete modules are promoted for reuse. In effect, we trade a small amount of upfront engineering for artifacts that are safer to compose and easier to review.
@@
Reproducibility further motivates our design. Community guidance stresses executable artifacts, clear dependency specifications, and explicit provenance as practical levers for robust claims [Pineau2021Reprod]. Accordingly, NN-RAG aligns reconstruction with Python’s import semantics and records dependency information using standard specifiers, so regeneration behaves predictably across machines and time [PyImportDocs, PEP508]. The outcome is a block library that supports faithful reruns, ablations, and fair comparisons without ad-hoc patching.
@@
Finally, NN-RAG is deliberately beyond code search. Classic clone detectors efficiently retrieve similar fragments, but they do not produce standalone, dependency-closed modules ready for drop-in reuse [deckard07, sajnani2016sourcerercc]. By lifting structural information (symbols, call graphs, import relationships) and closing the dependency graph before validation, our pipeline moves from “find something like this” to “provide a verified module that works here.” This enables fast, controlled experimentation and surfaces rarely co-deployed practices that, when recombined, can improve quality without inventing bespoke blocks. In the context of this paper, the same neutral specifications that back reliable reuse also furnish instruction–target pairs to fine-tune an LLM on unique ideas, allowing the model to propose or adapt designs without re-crawling the entire corpus at inference time.
@@ 3 Related Work
Curated libraries and model zoos have dramatically lowered adoption costs for known architectures by providing standardized implementations, training utilities, and checkpoints behind stable APIs. Representative efforts include timm for image backbones and training recipes, Detectron2 for detection/segmentation, and the OpenMMLab toolboxes such as MMDetection; PyTorch Hub and the Hugging Face Hub further generalize distribution and discovery across tasks. These ecosystems excel at consistency and breadth within a maintained repository or registry, but they do not attempt cross-repository discovery or automatic, dependency-closed assembly of new modules from heterogeneous sources—the specific gap NN-RAG targets. [timmdocs, detectron2git, mmdetectiongit, mmdetectionpaper, torchhubdocs, huggingfacehubdocs, torchvisionmodels]
@@
A separate line of work retrieves similar code rather than packaging executable modules. Classic clone detectors—DECKARD (tree-based), SourcererCC (token/index-based), and NiCad (near-miss, pretty-printing/normalization)—scale to large corpora and recover exact or near-miss clones with strong precision/recall. Deep code-search methods (e.g., DeepCS/CODEnn) go beyond surface similarity by learning joint embeddings of natural-language queries and code. These systems are effective for finding fragments but typically stop short of producing standalone, dependency-closed Python modules with import-preserving regeneration and runtime validation; bridging that last mile is the focus of our pipeline. [deckard07, sajnani2016sourcerercc, nicad09, deepcs18, deepcsdl]
@@
A fast-growing body of research studies repository-level generation and agentic editing. RepoCoder integrates iterative retrieval with a code LLM for repository-level completion and releases the RepoEval benchmark; more recent agent frameworks (e.g., SWE-agent) add explicit interfaces for editing files, running tests, and navigating projects. Benchmarks such as SWE-bench (and its Lite and “plus” variants) evaluate end-to-end issue resolution in real repositories, and follow-on systems (e.g., CodeR) explore multi-agent task graphs. This literature operationalizes how to read, modify, and evaluate full repos with LLMs. By contrast, NN-RAG prioritizes retrieval + assembly + validation to create dependency-closed, verified modules first, and only then (optionally) feeds their neutral specifications into LLM workflows; the two directions are complementary rather than competing. [repocoder23, repocoderpdf, sweagent24neurips, sweagentgit, swebench23, swebenchlite, swebenchplus24, coder24]
@@
Finally, we note that many production-quality vision stacks and registries (e.g., TorchVision’s model collection and the Hugging Face Hub) provide pre-trained artifacts with documented weights and APIs, which we leverage as downstream consumers once NN-RAG has produced executable modules; their goals (distribution and reproducibility) are orthogonal to NN-RAG ’s cross-repo extraction and validation. [torchvisionmodels, huggingfacehubdocs]
@@ 4.1 Repository Configuration
To contextualize this configuration, Figure 2 summarizes the distribution of extracted neural network blocks per repository. It presents a horizontal bar chart titled Neural network blocks by repository with the x-axis labeled Blocks extracted and a red dashed vertical line marking the overall mean ($\approx 162$ blocks). The largest share arises from the aggregated “Others (12 repos)” group (335 blocks), underscoring a long tail of useful but individually smaller projects beyond the marquee libraries. Among the named repositories, huggingface/transformers (198) and pytorch-image-models (167) exceed the mean, reflecting their breadth of reusable layers and pretrained models. pytorch/vision (156) lies just below the mean, consistent with its focused scope on canonical vision components. Detection-centric libraries (open-mmlab/mmdetection, 142; open-mmlab/mmsegmentation, 118; facebookresearch/detectron2, 89) and infrastructure/tooling (open-mmlab/mmcv, 94) contribute substantial but below-mean counts. Together with Table 1, this distribution indicates that our index covers both high-impact hubs and the broader ecosystem, enabling comprehensive retrieval across tasks and modalities.
@@ 4.3 Extraction Pipeline
Our end-to-end pipeline comprises seven phases as shown in Figure 3: (1) Automated block discovery—we statically scan configured repositories to identify class definitions that inherit from nn.Module, are non-abstract, and implement forward(), producing a JSON list of candidate names; (2) Repository cloning and caching—repositories are shallow-updated into a local cache to minimize network and disk overhead and to enable concurrent access by downstream phases [git_clone]; (3) LibCST-based parsing and indexing—each file is round-tripped through LibCST to retain concrete syntax while collecting symbol tables, imports, and module-level constants [libcst]; artifacts are persisted in SQLite with SHA-1 content digests for incremental rebuilds [sqlite_docs, python_sqlite3, python_hashlib]; (4) Symbol discovery and import graph—we register fully qualified symbols and construct a directed import graph for efficient dependency traversal, adhering to the semantics of Python’s import machinery [python_import_system]; (5) Scope-aware dependency resolution—free-name analysis respects the LEGB model (local, enclosing, global, built-in) and ranks candidates by resolution confidence (direct import, qualified name, heuristic match) [python_execution_model]; (6) Import-preserving code generation—resolved dependencies are emitted in definition-before-use order via topological sorting while preserving original import forms (including from X import Y as Z) [python_graphlib, python_import_system]; and (7) Validation and QA—generated modules are checked by AST parsing and bytecode compilation before sandboxed execution; successful builds are promoted to the registry and failures retain diagnostics for remediation [python_ast, python_compile]. Concretely, the resolver unifies static analysis with import-graph traversal: it computes a recursive transitive closure that handles cycles gracefully, implements LEGB semantics for comprehensions and assignment expressions, treats module constants and aliases as first-class dependencies, and preserves import statements verbatim to maintain runtime behavior [python_execution_model, python_import_system]. In practice, we parallelize parsing and extraction with concurrent.futures, keep a persistent SQLite store, order imports per PEP 8 [pep8], and compute cache keys with SHA-1 [python_hashlib].
@@
Interface and reproducibility.
For reproducibility, we expose a single CLI entry point and a minimal Python API. The CLI is invoked as python3 -m ab.rag (help via --help); common actions include extracting a single block (--block ResNet), extracting a small set (--blocks ResNet VGG DenseNet), or file-driven extraction using the default nn_block_names.json. Programmatically, BlockExtractor provides warm_index_once() to prepare clones and an index, then extract_single_block("ResNet") and extract_multiple_blocks([...]); a file-based variant supports limits and restart points [nnrag_repo].
@@
Our extractor never redistributes third-party source files. Instead, it indexes repositories to recover design information (signatures, invariants, dependency graphs) and emits neutral specifications used downstream. All original files remain in their upstream repos; our artifacts record repository identifiers and file hashes for traceability without copying code.
@@ 5 Experiments
Our evaluation (under the 19-repo configuration in Sec. 4.2) discovered 1,289 neural network block names. The blocks cover attention mechanisms, convolutional layers, transformer components, losses, pooling, normalization, and higher-level architectures; Table 2 summarizes the distribution. Consistent with the LEMUR corpus, our evaluation therefore prioritizes vision blocks and tasks, using a vision-centric repository configuration while keeping all mechanisms of NN-RAG unchanged
@@
Extraction configuration. We run the end-to-end pipeline with an indexing policy of missing (incremental re-indexing from a persistent SQLite cache [sqlite_docs, python_sqlite3]), dynamic parallelism $\max(\mathrm{CPU\_count},8)$ via concurrent.futures [python_concurrent_futures], and automatic validation enabled. Failed extractions use a two-attempt exponential backoff; cache validity is enforced with SHA-1 content digests [python_hashlib].
Evaluation metrics. We report: (i) extraction success rate, (ii) validation pass rate (AST parse, bytecode compilation, sandboxed execution [python_ast, python_compile]), (iii) dependency resolution accuracy, and (iv) a code quality score.
@@
Quality analysis. Manual inspection of 50 random blocks showed: (i) formatting preserved in 98%, (ii) dependency completeness in 94%, and (iii) PEP 8 import organization [pep8].
Performance and scalability. Cold-start indexing of 19 repositories completes in $\sim$5–10 minutes; subsequent runs under the missing policy complete in <30<30 seconds owing to persistent caching and content-hash lookups.
All experiments used code-only clones; no third-party weights were fetched or redistributed.

### v2

@@ Abstract
Reusing existing neural-network components is critical for research efficiency, yet identifying, extracting, and validating such modules across large open-source codebases remains challenging. We introduce NN-RAG, a retrieval-augmented generation system that transforms heterogeneous PyTorch repositories into a searchable, executable library of validated neural modules. Unlike conventional code search or clone-detection tools, NN-RAG performs scope-aware dependency resolution, import-preserving reconstruction, and validator-gated promotion, ensuring that retrieved modules are scope-closed and compilable. NN-RAG surpasses state-of-the-art approaches in extracting algorithmic logic from open-source repositories, achieving a 73.0% task success rate and 91.4% zero-shot executability—substantially outperforming general-purpose code agents lacking deterministic scope-aware protocols. Applied to 19 major repositories, the system extracted 1,289 candidate blocks, validated 941, and found that over 80% are structurally unique. Through multi-level de-duplication, NN-RAG contributes approximately 72% of all novel architectures in the LEMUR dataset. Beyond extraction, it enables automated cross-repository migration of architectural patterns, regenerating dependency-complete modules in new contexts at scale—a capability not provided by existing open-source systems. Overall, NN-RAG converts fragmented vision code into a reproducible, provenance-tracked substrate for algorithmic discovery. By formalizing the extraction of dependency-closed neural primitives, this work establishes a scalable methodology for ‘Neural Architecture Mining’, transforming disparate source code into a standardized, executable dataset for research.
Machine Learning, ICML
@@ 1 Introduction
We address this gap with NN-RAG, a retrieval-first framework that converts heterogeneous Python repositories into dependency-closed, executable modules with recorded provenance. The framework operationalizes a simple hypothesis: reliable retrieval, assembly, and validation at an appropriate granularity can both reduce the overhead of reuse (e.g., ablations and baseline construction) and expose architectural combinations that are underrepresented in the literature. In addition, NN-RAG produces neutral module specifications that can be used to fine-tune language models on reusable design principles while discouraging memorization of upstream implementations.
The goal of this paper is not to pursue leaderboard performance, but rather the discovery and extraction of new ideas. Nevertheless, while introducing new and unique ideas to the LEMUR dataset (Goodarzi et al., 2025), the pipeline surfaced a model that currently attains the best accuracy on that benchmark (Fig. 7). This outcome supports our goal: grounding design decisions in retrieved, validator-gated modules can improve both research velocity and model quality. To understand why, we complement headline numbers with targeted ablations, isolating which ingredients contribute most to the observed gains.
@@
Using a companion curation tool (NN-DUP (48)) with exact, MinHash/LSH near-deduplication, AST-fingerprint structural deduplication, and a diversity top-up, we find that the overwhelming majority of unique architectures in LEMUR originate from NN-RAG extractions. (A. T. Goodarzi, R. Kochnev, W. Khalid, F. Qin, T. A. Uzun, Y. S. Dhameliya, Y. K. Kathiriya, Z. A. Bentyn, D. Ignatov, and R. Timofte (2025); 49; BigCode (2023))
@@
Roadmap. Section 2 motivates the retrieval-first stance and summarizes findings; Section 3 explains the related work; Section 4 details the pipeline; Section 5 reports extraction/validation statistics and ablations; Section 6 dicusses the results; Section B outlines license compliance; Section 7 concludes.
@@ 2 Motivation
Although the community has produced a large body of reusable layers, losses, and architectural motifs, these contributions remain dispersed across heterogeneous PyTorch repositories with inconsistent conventions and nontrivial dependency structure. Consequently, adopting even a modest component typically requires (i) discovery at an appropriate granularity, (ii) closure of transitive dependencies consistent with Python import semantics, and (iii) validation of an executable artifact suitable for independent reproduction. This overhead reduces iteration speed, complicates fair empirical comparison, and limits the practical scientific value of released code.
@@
Our stance in NN-RAG is therefore retrieval first. Rather than ask a language model to synthesize code from scratch, we ground assembly in retrieved, verifiable sources, and only optionally invoke generation downstream from neutral specifications. Retrieval-augmented methods are known to improve factual grounding by supplementing a model’s parametric memory with explicit context; in our setting, that means the system privileges concrete, inspectable code over unconstrained paraphrase (Lewis et al., 2020; Huang et al., 2023). This design choice directly serves our aims: it turns the open-source corpus into a dependable stream of reusable ideas while curbing the brittleness and hallucination risks of pure free-form generation.
@@
Reliability and safety concerns reinforce this choice. Controlled studies show that AI coding assistance can encourage plausible but insecure solutions when specifications are implicit (Perry and others, 2023). NN-RAG counters this with validator-gated artifacts: import-preserving regeneration followed by static/dynamic checks, so only dependency-complete modules are promoted for reuse. In effect, we trade a small amount of upfront engineering for artifacts that are safer to compose and easier to review.
@@
Reproducibility further motivates our design. Community guidance stresses executable artifacts, clear dependency specifications, and explicit provenance as practical levers for robust claims (Pineau and others, 2021). Accordingly, NN-RAG aligns reconstruction with Python’s import semantics and records dependency information using standard specifiers, so regeneration behaves predictably across machines and time (Python Software Foundation (2025); 51). The outcome is a block library that supports faithful reruns, ablations, and fair comparisons without ad-hoc patching.
@@
Finally, NN-RAG is deliberately beyond code search. Classic clone detectors efficiently retrieve similar fragments, but they do not produce standalone, dependency-closed modules ready for drop-in reuse (Jiang et al., 2007; Sajnani et al., 2016). By lifting structural information (symbols, call graphs, import relationships) and closing the dependency graph before validation, our pipeline moves from “find something like this” to “provide a verified module that works here.” This enables fast, controlled experimentation and surfaces rarely co-deployed practices that, when recombined, can improve quality without inventing bespoke blocks. In the context of this paper, the same neutral specifications that back reliable reuse also furnish instruction–target pairs to fine-tune an LLM on unique ideas, allowing the model to propose or adapt designs without re-crawling the entire corpus at inference time.
@@ 3 Related Work
Curated libraries and model zoos have dramatically lowered adoption costs for known architectures by providing standardized implementations, training utilities, and checkpoints behind stable APIs. Representative efforts include timm for image backbones and training recipes, Detectron2 for detection/segmentation, and the OpenMMLab toolboxes such as MMDetection; PyTorch Hub and the Hugging Face Hub further generalize distribution and discovery across tasks. These ecosystems excel at consistency and breadth within a maintained repository or registry, but they do not attempt cross-repository discovery or automatic, dependency-closed assembly of new modules from heterogeneous sources—the specific gap NN-RAG targets (62; 15; 46; K. Chen, J. Huang, et al. (2019); 73; 33; 74).
@@
A separate line of work retrieves similar code rather than packaging executable modules. Classic clone detectors—DECKARD (tree-based), SourcererCC (token/index-based), and NiCad (near-miss, pretty-printing/normalization)—scale to large corpora and recover exact or near-miss clones with strong precision/recall. Deep code-search methods (e.g., DeepCS/CODEnn) go beyond surface similarity by learning joint embeddings of natural-language queries and code. These systems are effective for finding fragments but typically stop short of producing standalone, dependency-closed Python modules with import-preserving regeneration and runtime validation; bridging that last mile is the focus of our pipeline (Jiang et al., 2007; Sajnani et al., 2016; Roy and Cordy, 2009; Gu et al., 2018a, b).
@@
A fast-growing body of research studies repository-level generation and agentic editing. RepoCoder integrates iterative retrieval with a code LLM for repository-level completion and releases the RepoEval benchmark; more recent agent frameworks (e.g., SWE-agent) add explicit interfaces for editing files, running tests, and navigating projects. Benchmarks such as SWE-bench (and its Lite and “plus” variants) evaluate end-to-end issue resolution in real repositories, and follow-on systems (e.g., CodeR) explore multi-agent task graphs. This literature operationalizes how to read, modify, and evaluate full repos with LLMs. By contrast, NN-RAG prioritizes retrieval + assembly + validation to create dependency-closed, verified modules first, and only then (optionally) feeds their neutral specifications into LLM workflows; the two directions are complementary rather than competing (F. Zhang, B. Chen, Y. Zhang, J. Keung, J. Liu, D. Zan, Y. Mao, J. Lou, and W. Chen (2023); 63; J. Yang and et al. (2024); 70; C. E. Jimenez and et al. (2023); 71; R. Aleithan and et al. (2024); D. Chen and et al. (2024)).
@@
Finally, we note that many production-quality vision stacks and registries (e.g., TorchVision’s model collection and the Hugging Face Hub) provide pre-trained artifacts with documented weights and APIs, which we leverage as downstream consumers once NN-RAG has produced executable modules; their goals (distribution and reproducibility) are orthogonal to NN-RAG ’s cross-repo extraction and validation (74; 33).
@@ 4 Methodology
Building upon prior architectural synthesis work conducted within the NNGPT framework (Kochnev et al., 2025; Jesani et al., 2025; Duvvuri et al., 2026; Mittal et al., 2025; Shrestha et al., 2026; Khalid et al., 2026; Uzun et al., 2026a; Gu et al., 2026; Adhikari et al., 2026), and drawing on the LEMUR dataset of diverse high-capacity and edge-optimized neural network models (Goodarzi et al., 2025; Uzun et al., 2026b; Din et al., 2025; Kumar et al., 2026; Kayani et al., 2026), we introduce NN-RAG, a system developed as an integral component for the NNGPT framework (Kochnev et al., 2026).
@@ 4.1 Repository Configuration
Figure 2 summarizes the per-repository distribution of extracted blocks (full discussion in Appendix A).
@@ 4.3 Extraction Pipeline
Our end-to-end pipeline comprises seven phases as shown in Figure 3: (1) Automated block discovery—we statically scan configured repositories to identify class definitions that inherit from nn.Module, are non-abstract, and implement forward(), producing a JSON list of candidate names; (2) Repository cloning and caching—repositories are shallow-updated into a local cache to minimize network and disk overhead and to enable concurrent access by downstream phases (20); (3) LibCST-based parsing and indexing—each file is round-tripped through LibCST to retain concrete syntax while collecting symbol tables, imports, and module-level constants (Meta Open Source, 2024); artifacts are persisted in SQLite with SHA-1 content digests for incremental rebuilds (67; 69; 28); (4) Symbol discovery and import graph—we register fully qualified symbols and construct a directed import graph for efficient dependency traversal, adhering to the semantics of Python’s import machinery (Python Software Foundation, 2024f); (5) Scope-aware dependency resolution—free-name analysis respects the LEGB model (local, enclosing, global, built-in) and ranks candidates by resolution confidence (direct import, qualified name, heuristic match) (Python Software Foundation, 2024c); (6) Import-preserving code generation—resolved dependencies are emitted in definition-before-use order via topological sorting while preserving original import forms (including from X import Y as Z) (24; Python Software Foundation (2024f)); and (7) Validation and QA—generated modules are checked by AST parsing and bytecode compilation before sandboxed execution; successful builds are promoted to the registry and failures retain diagnostics for remediation (Python Software Foundation, 2024a, b). Concretely, the resolver unifies static analysis with import-graph traversal: it computes a recursive transitive closure that handles cycles gracefully, implements LEGB semantics for comprehensions and assignment expressions, treats module constants and aliases as first-class dependencies, and preserves import statements verbatim to maintain runtime behavior (Python Software Foundation, 2024c, f). In practice, we parallelize parsing and extraction with concurrent.futures, keep a persistent SQLite store, order imports per PEP 8 (van Rossum et al., 2025), and compute cache keys with SHA-1 (28).
@@
We provide a CLI and minimal Python API for reproducible extraction; usage details are in Appendix A.
@@
Our extractor never redistributes third-party source files. Instead, it indexes repositories to recover design information (signatures, invariants, dependency graphs) and emits neutral specifications used downstream. All original files remain in their upstream repos; our artifacts record repository identifiers and file hashes for traceability without copying code, See the Appendix B for license compliance details.
@@ 5 Experiments
Our evaluation (under the 19-repo configuration in Sec. 4.2) discovered 1,289 neural network block names. The blocks cover attention mechanisms, convolutional layers, transformer components, losses, pooling, normalization, advanced augmentation techniques (inspired by (Aboudeshish et al., 2025)), and higher-level architectures; Table 2 summarizes the distribution. Consistent with the LEMUR corpus, our evaluation therefore prioritizes vision blocks and tasks, using a vision-centric repository configuration while keeping all mechanisms of NN-RAG unchanged
@@
Extraction configuration. We run the end-to-end pipeline with an indexing policy of missing (incremental re-indexing from a persistent SQLite cache (67; 69)), dynamic parallelism $\max(\mathrm{CPU\_count},8)$ via concurrent.futures (13), and automatic validation enabled. Failed extractions use a two-attempt exponential backoff; cache validity is enforced with SHA-1 content digests (28).
Evaluation metrics. We report: (i) extraction success rate, (ii) validation pass rate (AST parse, bytecode compilation, sandboxed execution (Python Software Foundation, 2024a, b)), (iii) dependency resolution accuracy, and (iv) a code quality score.
@@
Quality analysis. Manual inspection of 50 random blocks showed: (i) formatting preserved in 98%, (ii) dependency completeness in 94%, and (iii) PEP 8 import organization (van Rossum et al., 2025).
Runtime and scalability characteristics are reported in Appendix C.
@@
5.1 Comparative Analysis of Extraction Efficacy
To evaluate the necessity of scope-aware extraction, we compare NN-RAG against four state-of-the-art baselines for code manipulation and ML automation. As shown in Table 3, while generalist agents demonstrate proficiency in localized edits, they do not optimize for the structural integrity required for modular neural network execution.
Table 3: Comparative analysis of module extraction efficacy. While general-purpose benchmarks (SWE-bench, GitTaskBench) evaluate broad issue resolution, NN-RAG is specialized for extracting dependency-closed neural primitives. ✓and × denote native support for structural verification; — indicates metrics not reported in the respective original literature.
Methodology NAS Task Success Executable Dependency Fully Scope
Targeted Rate (%) Ratio (%) Closure (%) Auto. Aware
SWE-bench (Yang et al., 2025) × 13.8 × ×
GitTaskBench (Ni et al., 2025) × 48.2 × ×
Manas (Venkatesh et al., 2024) ×
Paper2Code (Going-Doer et al., 2025) ×
NN-RAG (Ours) 73.0 91.4 98.7
To assess the efficacy of scope-aware extraction, we evaluate NN-RAG against four state-of-the-art baselines representing distinct paradigms in code manipulation and ML automation: (1) issue-centric agents (SWE-bench (Yang et al., 2025)), (2) task-oriented agents (GitTaskBench (Ni et al., 2025)), (3) mining-based AutoML (Manas (Venkatesh et al., 2024)), and (4) literature-to-code synthesis (Paper2Code (Going-Doer et al., 2025)). As illustrated in Table 3, while generalist agents demonstrate proficiency in localized logic edits or repository-scale task resolution, they typically fail to maintain the structural integrity essential for modular neural network execution.
Agents evaluated on SWE-bench and GitTaskBench are primarily optimized for defect correction and feature implementation; consequently, they exhibit limited zero-shot executability when isolating standalone modules, often due to unresolved global dependencies. While specialized frameworks like Manas leverage repository mining for model transfer, they lack a deterministic scope analysis for high-precision sub-module extraction. Similarly, although Paper2Code incorporates architectural planning for de novo synthesis, it is not architected to extract internal logic from existing heterogeneous source code. In contrast, NN-RAG employs a specialized AST-rewriting protocol to ensure near-total dependency closure (98.7%). By integrating native scope awareness with a gated validation system, our approach achieves superior structural reliability and a significantly higher task success rate (73.0%) within the neural architecture domain.